Evaluate the value of the finite population correction factor when a sample of size 10
drawn from the population having size 100.
1
Expert's answer
2020-08-25T15:08:11-0400

Formula for finite population correction factor is given by,


"fpc = \sqrt{\frac{N-n}{N-1}}"


where,

sample size n = 10

population size N = 100


Hence,

"fpc = \sqrt{\frac{100-10}{100-1}} = \sqrt\frac{90}{99} = 0.95346"